OpenStudy (anonymous):

Evaluate the expression when x = –24. x-12/6 A –6 B. -2 C. 2 D. 6

OpenStudy (anonymous):

(-24-12)/6 = -36/6=-6, so A

OpenStudy (kittiwitti1):

\[f(x)= \frac{ x-12 }{ 6 }\]if x = -24\[f(-24) = \frac{ -24-12 }{ 6 } = \frac{ -48 }{ 6 } = -8\]

OpenStudy (kittiwitti1):

You have to set up the entire equation to get full credit for most teachers, you know. @beans1992

OpenStudy (anonymous):

-24-12 = -36, not -48

OpenStudy (anonymous):

Thank you!!

OpenStudy (kittiwitti1):

Oh yeah. Whoops sorry ! Thanks for correcting @beans1992